g(f(2))
f(x)=-3x+2,g(x)=x/5

Answer by ewatrrr(24785)


Hi,
f(x)=-3x+2
f(2) = -3*2 + 2 = -4
g(x)=x/5
g(f(2))= g(-4) = -4/5